The calculation of a weighted average cost of capital (WACC) involves calculating the weighted average of the required rates of return on debt and equity, where the weights equal the percentage of each type of financing in the firm’s overall capital structure.

1.) __________ is the symbol that represents the cost of raising capital through retained earnings in the weighted average cost of capital (WACC) equation.

2.) $Tim Co. has 1.13 million of debt, $1.6 million of preferred stock, and $3.04 million of common equity. The appropriate weight of the firm's common equity debt in the calculation of the company's weighted average cost of capital is __________ (19.58%, 22.18%, 30.50%, 52.69%)
Answer #1

1)

The symbol for cost of retained earnings is 'rs'

2)

value of firm = 1.13 + 1.6 + 3.04

= 5.77 million

weight of debt = 1.13/5.77

= 19.58%

weight of equity = 3.04/5.77

= 52.69%

Weight of preferred stock = 1.6/5.77

= 27.73%